# pnasresearcharticle

### v2

@@ Abstract
The epidemiological dynamics of Mycoplasma pneumoniae are characterized by complex and poorly understood multiannual cycles, posing challenges for forecasting. Using Bayesian methods to fit a seasonally forced transmission model to long-term surveillance data from Denmark (1958–1995, 2010–2025), we investigate the mechanisms driving recurrent outbreaks of M. pneumoniae. The period of the multiannual cycles (predominantly approx. 5 years in Denmark) are explained as a consequence of the interaction of two time-scales in the system, one intrinsic and one extrinsic (seasonal). While it provides an excellent fit to shorter time series (a few decades), we find that the deterministic model eventually settles into an annual cycle, failing to reproduce the observed 4-5-year periodicity long-term. Upon further analysis, the system is found to exhibit transient chaos and thus high sensitivity to stochasticity. We show that environmental (but not purely demographic) stochasticity can sustain the multi-year cycles via stochastic resonance. The disruptive effects of COVID-19 non-pharmaceutical interventions (NPIs) on M. pneumoniae circulation constitute a natural experiment on the effects of large perturbations. Consequently, the effects of NPIs are included in the model and medium-term predictions are explored. Our findings highlight the intrinsic sensitivity of M. pneumoniae dynamics to perturbations and interventions, underscoring the limitations of deterministic epidemic models for long-term prediction. More generally, our results emphasize the potential role of stochasticity as a driver of complex cycles across endemic and recurring pathogens.
@@
Mycoplasma pneumoniae |||| Epidemiological dynamics |||| Dynamical systems |||| Chaos |||| Stochasticity
@@
This manuscript was compiled on April 16, 2025
@@
The bacterial pathogen Mycoplasma pneumoniae is simultaneously common and unusual. It lacks a cell wall and is among the smallest free-living bacteria (1). The infection it causes has been given names such as (primary) atypical pneumonia (2), cold pneumonia and walking pneumonia (3) due to its tendency to cause less severe disease compared with typical bacterial pneumonia. It is a very common respiratory pathogen, especially in children and adolescents, and accounts for 15-20% of all cases of pneumonia (4). Disease courses are typically mild, but severe outcomes do occur. About 5-10% of upper airway infections with M. pneumoniae progress to pneumonia, and there are likely many asymptomatic cases (5), suggesting a high degree of under-ascertainment. Extrapulmonary symptoms are increasingly recognized (6, 7) and include complications such as encephalitis and RIME/Stevens-Johnson Syndrome (8, 9), with M. pneumoniae being the most common infectious cause of the latter in children (10, 11). As many as 25% of M. pneumoniae cases develop some extrapulmonary symptoms (12), most commonly dermatological or gastrointestinal; sometimes cardiovascular (13) and neurological (14) complications are seen, the latter in as many as 1-10% of those M. pneumoniae cases which require hospitalization (15, 12, 16).
@@
From an epidemiological dynamics perspective, a remarkable feature of M. pneumoniae are the observed complex outbreak cycles, most often with a 3-7 year period (17, 18, 19, 20). In Denmark (2), as well as England and Wales (21), 4-5 year cycles have generally been observed. This is in contrast to some classical, strongly immunizing childhood diseases, such as measles, which tend to exhibit simpler annual or biennial dynamics (22), although there are exceptions to that rule, e.g. rubella (23, 24). Despite being such a common pathogen, the mechanisms behind the complex epidemiological dynamics of M. pneumoniae remain poorly understood, complicating prediction and forecasting. In general, predictability is important for timely resource allocation and deployment of relevant screening and diagnostic testing schemes. When the year-to-year incidence exhibits wide variability, planning for outbreaks becomes simultaneously more important, and more difficult. Discerning and understanding patterns in this variability is thus paramount for effective planning, not least for M. pneumoniae, where severe outcomes are typically preventable if diagnosed and treated in a timely manner.
@@
Since early 2020, the multiannual cycles of M. pneumoniae have been largely disrupted (25, 26). This is an example of the polymicrobial effects of the COVID-19-related non-pharmaceutical interventions (NPIs), observed as a disruption of circulation across many pathogens (27) but perhaps evidenced most strikingly by the apparent extinction of the Yamagata strain of influenza B virus (28, 29).
@@
Over three years, M. pneumoniae circulation was drastically curtailed in many locations and Denmark is no exception. A similar multi-year delay in M. pneumoniae resurgence has recently been reported for Europe and Asia (30) as well as the United States (25, 31). Such large-scale perturbations effectively provide a “natural experiment”, demonstrating how the dynamics of M. pneumoniae responds to a prolonged disruption and the associated build-up of susceptible individuals, as well as aiding model parameter estimation.
@@
Here, we use a mathematical model that incorporates seasonality, vital dynamics (births/deaths), immunity loss and the effects of NPIs to study recent (2010–2025) and historical (1958–1995) M. pneumoniae incidence data from Denmark. The simple deterministic model is found to capture the observed dynamics well over a few decades but eventually reverts to annual outbreaks, failing to reproduce the prominent multiannual cycles long-term. We show that the presence of environmental noise can indefinitely sustain the multi-year outbreak patterns. In doing so, we highlight the important role that fluctuations in contact patterns, environmental noise, and demographic processes plays when forecasting pathogens with complex dynamics such as M. pneumoniae.
@@
We utilize two different data sets as correlates of M. pneumoniae incidence, both from Denmark (Statens Serum Institut, SSI). One has M. pneumoniae-positive PCR tests (and the raw number of tests performed) with a weekly reporting frequency and covers the period 2010-2025. This data set forms the primary basis for parameter estimation, since the test data is highly M. pneumoniae-specific and temporally well-resolved. Secondarily, we use a data set covering the period 1958-1995. Beginning in 1946, patient blood samples collected at hospitals and general practitioners for the diagnosis of Primary Atypical Pneumonia (later identified as M. pneumoniae) were sent to SSI to undergo the cold agglutinin (CA) reaction. This reaction, unfortunately, has only low to moderate sensitivity ($\sim 20-80\%$ (32, 2)) and moderate specificity ($\sim 82-92\%$ (32, 2)) for M. pneumoniae. However, beginning in 1958, CA-positive samples were also subjected to M. pneumoniae Complement Fixation (MpCF), which was found to have higher specificity (33, 34). We thus also treat those tests from the 1958-1995 data set which are simultaneously CA- and MpCF-positive as an incidence proxy on which we base (secondary) parameter estimates.
@@
The duration of immunity following M. pneumoniae infection is known not to be life-long, but the precise duration is unknown, with estimates ranging from a few years to more than a decade (35, 2, 25). For this reason, we opted for an SIRS (Susceptible-Infected-Recovered-Susceptible) compartmental model which allows previously infected individuals to become susceptible, and thus potentially reinfected. The model includes seasonal modulation of the transmission rate and equal birth/death rates (which are assumed constant within each data fit, for simplicity). The duration of immunity was treated as a parameter to be estimated during fitting to data.
@@
The influence of NPIs are implemented as a slowly-varying multiplicative effect on the transmission rate from March 2020 onward, representing a reduction in overall contact rates, the magnitude of which is an estimated parameter (see Supplementary Fig. S5).
@@
We estimate a basic reproductive number of $R_{0}=1.8$ (95%CI [1.7, 2.0], in good agreement with ref. (36)), a mean duration of immunity of 11 years and a relative amplitude of seasonality in transmission of $0.19$. All estimated parameters and their uncertainties can be found in Table S1.
@@
With a sinusoidal seasonality function [^1] We opted for a sinusoidal seasonality function for simplicity, but have also directly estimated the shape of the seasonality function from the data. The resulting seasonality function can be found in Supplementary Fig. S4., the model fits the 2010-2025 data well (with the partial exception of the 2017 season in which the apparent incidence was somewhat higher than modeled), see Fig. 1A. The historical 1958-1995 data, which is overall of much lower fidelity, is captured more loosely by the model, see Fig. 1B. Interestingly, here we had to allow for a slow change in the mean transmission rate and ascertainment (allowed to change only slightly every 4 years), otherwise the model would tend to settle into an annual pattern.
@@
This leads us to a more general observation: the SIRS model with M. pneumoniae parameters is able to capture the patterns reasonably well over a few decades, but in the absence of perturbations, the modeled dynamics always decays into an annual pattern eventually. This is of course at odds with the ubiquity of complex multiannual cycles in M. pneumoniae (2, 17, 18, 19, 20, 21) and other SIRS-like epidemiological systems such as parvovirus B19 (37, 38, 39, 40, 41). In the next section we discuss how the time-scales of the M. pneumoniae system may give rise to an approximately 5-year cycle, at least transiently. We then go on to investigate the role that stochasticity plays in sustaining such cycles over decades and in re-igniting them even after periods of approximately annual dynamics. In the final part of the paper, we study projections of future M. pneumoniae incidence with and without stochasticity.
Figure 1: Model fits to Mycoplasma pneumoniae incidence data. A seasonal SIRS model is fitted to recent (2010-2025) and historical (1958-1995) M. pneumoniae testing data from Denmark using Bayesian inference. A) The seasonal SIRS model explains observed M. pneumoniae dynamics 2010-2025 well. The fitted data are PCR positivity rates. B) The same type of model captures key features of observed M. pneumoniae dynamics 1958-1995. The fitted data are variance-stabilized numbers of samples which are simultaneously CA and MpCF-positive. (Data source: Statens Serum Institut and (2)).
Two time-scales conspire to create complex cycles
@@
The seasonally forced SIRS system with vital dynamics contains two comparable oscillatory time-scales: the one-year seasonal period and the time-scale set by the intrinsic epidemiological dynamics. To ascertain the latter time-scale, the corresponding autonomous system may be studied (that is, the SIRS system with vital dynamics but no seasonal forcing). This system will eventually reach an endemic equilibrium, but will oscillate as it approaches this equilibrium. A linearization of the autonomous system around its endemic equilibrium yields an intrinsic oscillation frequency as well as an exponential decay rate (42), given by a combination of the transmission rate $\beta$, immune waning rate $\delta$, recovery rate $\gamma$ and birth/death rate $\mu$. Using the parameter values estimated from the Bayesian fit to Danish data (2010-2025), the frequency is found to be $f_{1}=1/(43.46\text{wk})$ (95%CI: [1/44.0, 1/43.0]). This natural oscillatory time-scale offers a clue to how the 4-5-year cycles come about: denoting the seasonal frequency by $f_{0}=1/\text{yr}$ and the intrinsic frequency by $f_{1}$ as above, the combined effect of the two time-scales can be qualitatively understood by considering the product of two sinusoidal signals with different frequency:
$\sin(f_{0}t)\sin(f_{1}t)=\tfrac{1}{2}\left(\cos((f_{0}-f_{1})t)+\cos((f_{0}+f_{1})t)\right)$
the resulting signal thus contains a time-scale which is slower than either of the two component time-scales. In our case, the frequency $f_{0}-f_{1}$ corresponds to a period of 5.1 years (95%CI: [4.8, 5.5]). It should be noted that $5\text{yr}/43.46\text{wk}=6$ to within 1%, while 5 years is of course also an integer multiple of the period of the seasonal driver (1 year). This suggests that the system displays subharmonic resonance (the second mechanism for population cycling described in (43) and a well-known phenomenon in non-linear oscillatory systems (44, 45)).
@@
Grossman remarked (46, 42) that the SIR equilibrium (w. vital dynamics) tends to be only weakly stable, since the real part of the eigenvalue (the decay rate) is much smaller than the imaginary part (the frequency). We find the same phenomenon for the SIRS system with M. pneumoniae parameters, with the ratio $|\text{Im}\{\lambda\}/\text{Re}\{\lambda\}|\approx 14$ (with $\lambda$ being the relevant eigenvalue of the Jacobian matrix for the linearized system).
@@
A simple Fourier transform of the 2010-2020 (pre-COVID) data yields a dominant period of 5.1 years, while a Lomb-Scargle (LS) periodogram (47), which has finer frequency resolution, places the peak at 4.9 years. For the MpCF-based 1958-1995 time series, the LS periodogram shows the annnual signal as being the strongest, which makes sense given the observation of (2) that the multiannual outbreak pattern was somewhat overtaken by annual dynamics during the late 1970’s and early 1980’s. However, the strongest multiannual peak for the whole period (1958-1995) is at 4.1yrs (LS)/4.2yrs (Fourier). If the 1958-1975 period is viewed in isolation, the strongest peak (eclipsing the annual signal as well) is at 4.6yrs (LS)/4.3yrs (Fourier). See Supplementary Fig. S2 for Fourier and LS spectra of the incidence time series.
@@
Spurred by the observation that the deterministic model will always eventually decay into annual dynamics, it is natural to ask how a departure from determinism affects the dynamics. Demographic stochasticity is a simple and ubiquitous source of noise in population dynamics (including epidemiological dynamics) (48, 49). This type of stochasticity occurs in finite populations and is fundamentally due to
1) Individuals being discrete (as opposed to the real numbers $S$, $I$ and $R$ of a typical ordinary differential equation-based model), and
2) The presence of probability rates such as $\mu$, $\beta$, $\gamma$ and $\delta$, which implies stochasticity in the timing of birth/death, infection, recovery and immune waning in finite populations.
Both of these effects can be taken into account by stochastic simulation, as may be implemented using the Doob-Gillespie algorithm (details in Materials and Methods) (50). With such a seasonally-forced stochastic SIRS model with vital dynamics, we find that multi-year cycles may indeed be sustained by demographic noise, suggesting that the system exhibits a stochastic resonance (42). However, clear multiannual cycles can be sustained by demographic stochasticity only in populations of less than approximately 2 million individuals, see Fig. 2. Simple demographic stochasticity alone is thus unlikely to be the source of sustained multiannual cycles in M. pneumoniae.
Figure 2: Demographic stochasticity excites multi-year cycles in smaller populations. Doob-Gillespie simulations of a seasonally forced SIRS model parametrized with Mycoplasma pneumoniae parameters as determined by Bayesian fit to 2010-2025 data. A) In small populations (initial population size $N_{0}=20,000$), demographic stochasticity leads to noisy multi-year cycles. B) In intermediate populations ($N_{0}=250,000$), multi-year cycles are prominent and fairly regular. C) In larger populations ($N_{0}=2\times 10^{6}$), annual dynamics dominates as the impact of demographic stochasticity is smaller.
@@
We go on to consider the effects of extra-demographic or environmental stochasticity. Random fluctuations in the transmission rate $\beta(t)$ will be the primary source of environmental stochasticity considered, but we will also explore the effects of more generic exogenous shocks to the system.
@@
As shown in Fig. 3A, random fluctuations in the transmission rate tends to stimulate approximately 5-year cycles for a wide range of moderate noise levels. When stochasticity is very slight, the annual cycle dominates, and when environmental fluctuations are very strong, the epidemic dynamics lack predictable patterns and become dominated by large but rare outbreaks, appearing as a Fourier spectrum with all frequencies represented and no discernible peaks. Details of the implementation of stochastic transmission rates, which we base on (51), can be found in Materials and Methods. It should be noted that even when the annual peak is the highest, a subdominant peak with e.g. 50% of the peak power of the annual signal will lead to a prevalence timeseries that appears very clearly multiannual. The averaging process of Fig. 3A leads to broader and flatter multiannual peaks than those found in individual spectra, and for this reason Fig. 3B shows the spectra of a few individual realizations in which the peaks at approx. 5 years stand out.
@@
While the 5-year cycle is the most typical multiannual cycle at moderate noise levels (given M. pneumoniae parameters), other cycles are excited as well, as shown in Fig. 3, with approximately three-year as well as 6-7 year cycles being somewhat common at moderate-to-high noise levels ($\sigma>0.2$).

### v3

@@ Abstract
The epidemiological dynamics of Mycoplasma pneumoniae is characterized by poorly understood complex multiannual cycles. The origins of these cycles have long been debated, and multiple explanations of varying complexity have been suggested. Using Bayesian methods, we fit a dynamical model to half a century of M. pneumoniae surveillance data from Denmark (1958–1995, 2010–2025) and uncover a parsimonious explanation for the persistent cycles, based on the theory of quasicycles. The period of the multiannual cycle (approx. 5 years in Denmark) is explained by susceptible replenishment due, primarily, to loss of immunity. While an excellent fit to shorter time series (a few decades), the deterministic model eventually settles into an annual cycle, unable to reproduce the persistent cycles. We find that environmental stochasticity (e.g., varying contact rates) stabilizes the multiannual cycles and so does demographic noise, at least in smaller or incompletely mixing populations. The temporary disappearance of cycles during 1979–1985 is explained as a consequence of stochastic mode-hopping. The circulation of M. pneumoniae was recently disrupted by COVID-19 non-pharmaceutical interventions (NPIs), providing a natural experiment on the effects of large perturbations. Consequently, the effects of NPIs are included in the model and medium-term predictions are explored. Our findings highlight the intrinsic sensitivity of M. pneumoniae dynamics to perturbations and interventions, underscoring the limitations for long-term prediction. More generally, our findings provide further evidence for the role of stochasticity as a driver of complex cycles across endemic and recurring pathogens.
@@
Mycoplasma pneumoniae $|$ Epidemiological dynamics $|$ Dynamical systems $|$ Quasicycles $|$ Stochasticity
@@
This manuscript was compiled on November 25, 2025
@@
The bacterial pathogen Mycoplasma pneumoniae is simultaneously common and unusual. It lacks a cell wall and is among the smallest free-living bacteria (1). The infection it causes has been given names such as (primary) atypical pneumonia (2), cold pneumonia, and walking pneumonia (3) due to its tendency to cause less severe disease compared with typical bacterial pneumonia. It is a very common respiratory pathogen, especially in children and adolescents, accounting for 15–20% of all cases of pneumonia (4). Disease courses are typically mild, but severe outcomes do occur. About 5–10% of upper airway infections with M. pneumoniae progress to pneumonia, and there are likely many asymptomatic cases (5), suggesting a high degree of under-ascertainment. Extrapulmonary symptoms are increasingly recognized (6, 7) and include complications such as encephalitis and RIME/Stevens-Johnson Syndrome (8, 9), with M. pneumoniae being the most common infectious cause of the latter in children (10, 11). As many as 25% of M. pneumoniae cases develop some extrapulmonary symptoms (12), most commonly dermatological or gastrointestinal. Sometimes cardiovascular (13) and neurological (14) complications are seen, the latter in as many as 1–10% of those M. pneumoniae cases which require hospitalization (15, 12, 16).
@@
From an epidemiological dynamics perspective, a remarkable feature of M. pneumoniae are the observed complex outbreak cycles which generally exhibit a 3-7 year period (17, 18, 19, 20). In Denmark (2), as well as England and Wales (21), 4–5 year cycles have generally been observed. This is in contrast to some classical, strongly immunizing childhood diseases, such as measles, which tend to exhibit simpler annual or biennial dynamics (22). However, there are exceptions to that rule, for example rubella (23, 24). Despite being such a common pathogen, the mechanisms behind the complex epidemiological dynamics of M. pneumoniae remain poorly understood, complicating prediction and forecasting. In general, even short and medium term predictability are important for timely resource allocation and deployment of relevant screening and diagnostic testing schemes. However, when the year-to-year incidence exhibits wide variability, planning for outbreaks becomes simultaneously more important and more difficult. Discerning and understanding patterns in this variability is thus paramount for effective planning, not least for M. pneumoniae, where severe outcomes are typically preventable if diagnosed and treated in a timely manner.
@@
Since early 2020, the multiannual cycles of M. pneumoniae have been largely disrupted (25, 26). Such polymicrobial effects of the COVID-19-related non-pharmaceutical interventions (NPIs) have been observed across many pathogens (27). The most striking example is perhaps the apparent extinction of the Yamagata strain of influenza B virus (28, 29).
@@
Over three years, M. pneumoniae circulation was drastically curtailed in many locations, and Denmark is no exception. A similar multi-year delay in M. pneumoniae resurgence has recently been reported for Europe and Asia (30) as well as the United States (25, 31). Such large-scale perturbations effectively provide a “natural experiment”, demonstrating how the dynamics of M. pneumoniae responds to a prolonged disruption and the associated build-up of susceptible individuals. A recent study of respiratory syncytial virus (RSV) and human metapneumovirus (hMPV) found that including the NPIs the and post-NPI rebound could aid model parameter estimation (32), since including data from outside the regular, cyclic regime tends to improve identifiability.
While the origins of the multiannual cycles of M. pneumoniae are generally regarded as unclear (33, 21), a number of different mechanisms of varying complexity have been proposed in the literature. Among the earlier ones was the suggestion by (2) that immunity following infection simply has approximately the same duration as the observed cycle. This explanation, however, has difficulty accounting for how cycle periods can differ somewhat between countries. Without accounting for susceptible dynamics in detail, it is also unclear why the immunity duration should have to match the cycle period. Zhang et al. (34) suggested a combination of strain dynamics, contact network structure and waning immunity as the explanation, while Omori et al. (35) explored three different possible explanations: school-term forcing, a low-dispersion distribution of immunity duration, and interference between M. pneumoniae serotypes. They find that, of these three explanations, a less variable duration of immunity (compared with the typically assumed exponential waning) best captures the generally observed patterns in that it allows for incidence oscillations. Recently, the authors of (26) used a TSIR model to describe the post-COVID rebound of M. pneumoniae in Denmark as well as four UN regions (Europe, Asia, the Americas, and Oceania) but found that they required at least a 90% reduction in the transmission rate due to non-pharmaceutical interventions in order to fit the data.
@@
Here, we use a mathematical model that incorporates seasonality, vital dynamics (births/deaths), immunity loss and the effects of NPIs to study recent (2010–2025) and historical (1958–1995) M. pneumoniae incidence data from Denmark. The simple deterministic model is found to capture the observed dynamics well over a few decades but eventually reverts to annual outbreaks, failing to reproduce the prominent multiannual cycles long-term. We show that the presence of environmental noise can indefinitely sustain the multi-year outbreak patterns. In doing so, we highlight the important role that fluctuations in contact patterns, environmental noise, and demographic processes play when forecasting pathogens with complex dynamics such as M. pneumoniae.
@@
We utilize two different data sets as correlates of M. pneumoniae incidence, both from Denmark (Statens Serum Institut, SSI). One has M. pneumoniae-positive PCR tests (and the raw number of tests performed) with a weekly reporting frequency and covers the period 2010–2025. This data set forms the primary basis for parameter estimation, since the test data is highly M. pneumoniae-specific and temporally well-resolved. Secondarily, we use a data set covering the period 1958–1995. Beginning in 1946, patient blood samples collected at hospitals and general practitioners for the diagnosis of Primary Atypical Pneumonia (later identified as M. pneumoniae) were sent to SSI to undergo the cold agglutinin (CA) reaction. This reaction, unfortunately, has only low to moderate sensitivity ($\sim 20-80\%$ (36, 2)) and moderate specificity ($\sim 82-92\%$ (36, 2)) for M. pneumoniae. However, beginning in 1958, CA-positive samples were also subjected to M. pneumoniae Complement Fixation (MpCF), which was found to have higher specificity (37, 38). We thus also treat those tests from the 1958–1995 data set which are simultaneously CA- and MpCF-positive as an incidence proxy on which we base (secondary) parameter estimates.
@@
The duration of immunity following M. pneumoniae infection is known not to be life-long, but the precise duration is unknown, with estimates ranging from a few years to more than a decade (39, 2, 25). For this reason, we opted for an SIRS (Susceptible-Infected-Recovered-Susceptible) compartmental model which allows previously infected individuals to become susceptible, and thus potentially reinfected. The model includes seasonal modulation of the transmission rate and equal birth/death rates (which are assumed constant within each data fit, for simplicity). The duration of immunity was treated as a parameter to be estimated during fitting to data.
@@
The influence of NPIs is implemented as a slowly-varying multiplicative effect on the transmission rate from March 2020 onward, representing a reduction in overall contact rates, the magnitude of which is an estimated parameter (see Supplementary Fig. S7).
@@
We estimate a basic reproductive number of $R_{0}=1.6$ (95%CI [1.45, 1.82]), in good agreement with ref. (40), a mean duration of immunity of 8–9 years and a relative amplitude of seasonality in transmission of $0.18$. All estimated parameters and their uncertainties can be found in Table 1.
@@
With a sinusoidal seasonality function [^1] We opted for a sinusoidal seasonality function for simplicity, but have also directly estimated the shape of the seasonality function from the data. The resulting seasonality function can be found in Supplementary Fig. S5., the model fits the 2010–2025 data well (with the partial exception of the 2017 season in which the apparent incidence was somewhat higher than modeled), see Fig. 1A. The historical 1958–1995 data, which is overall of much lower fidelity, is captured more loosely by the model, see Fig. 1B. Interestingly, here we had to allow for a slow change in the mean transmission rate and ascertainment (allowed to change only slightly every 4 years), otherwise the model would tend to settle into an annual pattern.
@@
This leads us to a more general observation: the SIRS model with M. pneumoniae parameters is able to capture the patterns reasonably well over a few decades, but in the absence of perturbations, the modeled dynamics always decays into an annual pattern eventually. This is of course at odds with the ubiquity of complex multiannual cycles in M. pneumoniae (2, 17, 18, 19, 20, 21) and other SIRS-like epidemiological systems such as parvovirus B19 (41, 42, 43, 44, 45). In the next section we discuss how the time-scales of the M. pneumoniae system may give rise to an approximately 5-year cycle, at least transiently. We then go on to investigate the role that stochasticity plays in sustaining such cycles over decades and in re-igniting them even after periods of approximately annual dynamics. In the final part of the paper, we fit a stochastic model to the data and study projections of future M. pneumoniae incidence with and without stochasticity.
The origins of intrinsic cycles
@@
The seasonally forced SIRS system with vital dynamics contains two comparable oscillatory time-scales: the one-year seasonal period and the time-scale set by the intrinsic epidemiological dynamics. To ascertain the latter time-scale, the corresponding autonomous system may be studied (that is, the SIRS system with vital dynamics but no seasonal forcing). This system will eventually reach an endemic equilibrium, but will perform underdamped oscillations as it approaches this equilibrium. A linearization of the autonomous system around its endemic equilibrium yields an intrinsic oscillation frequency as well as an exponential decay rate (46), given by a combination of the transmission rate $\beta$, immune waning rate $\delta$, recovery rate $\gamma$ and birth/death rate $\mu$. See Eq. 5 in Methods and Materials for the analytic expression for the intrinsic period. Using the parameter values estimated from the Bayesian fit to Danish data (2010–2025), the intrinsic period is found to be $T=5.0\text{yr}$ (95%CI: [4.96, 5.10]). In Supplementary Fig. S6, we explore the pairwise parameter dependence of the intrinsic period. The parameter values estimated by fitting to the 2010–2025 data are indicated with black crosses. The intrinsic period depends strongly on the transmission rate ($\beta$), infectious period ($1/\gamma$) and immunity duration ($1/\delta$), while the birth/death rate ($\mu$) plays only a minor role. Since the mean duration of immunity and infectious period are primarily a function of intrinsic host-pathogen interactions, we conclude that the primary determinant of periodicity in different countries is likely to be the transmission rate.
@@
Grossman remarked (47, 46) that the SIR equilibrium (w. vital dynamics) tends to be only weakly stable, since the real part of the eigenvalue (the decay rate) is much smaller than the imaginary part (the frequency). We find the same phenomenon for the SIRS system with M. pneumoniae parameters, with the ratio $|\text{Im}\{\lambda\}/\text{Re}\{\lambda\}|\approx 12$ (with $\lambda$ being the relevant eigenvalue of the Jacobian matrix for the linearized system).
@@
A simple Fourier transform of the 2010–2020 (pre-COVID) data yields a dominant period of 5.1 years, while a Lomb-Scargle (LS) periodogram (48), which has finer frequency resolution, places the peak at 4.9 years. For the MpCF-based 1958–1995 time series, the LS periodogram shows the annual signal as being the strongest, which makes sense given the observation of (2) that the multiannual outbreak pattern was somewhat overtaken by annual dynamics during the late 1970’s and early 1980’s. However, the strongest multiannual peak for the whole period (1958–1995) is at 4.1yrs (LS)/4.2yrs (Fourier). If the 1958–1975 period is viewed in isolation, the strongest peak (eclipsing the annual signal as well) is at 4.6yrs (LS)/4.3yrs (Fourier). See Supplementary Fig. S2 for Fourier and LS spectra of the incidence time series, as well as a wavelet analysis which shows changes in the dominant frequency over time.
In the next two sections, we consider the role that stochasticity may play in sustaining the observed multiannual cycles. Apart from stochasticity being a fact of nature, the introduction of a stochastic model is further motivated by the observation that the fit in Fig. 1A displays low uncertainty compared with the variability in the observed data, suggesting that noise is needed to explain the data.
@@
Spurred by the observation that the deterministic model will always eventually decay into annual dynamics, it is natural to ask how a departure from determinism affects the dynamics. Demographic stochasticity is a ubiquitous source of noise in population dynamics (including epidemiological dynamics) (49, 50). This type of stochasticity occurs in finite populations and is fundamentally due to two factors. 1) Individuals being discrete (as opposed to the real numbers $S$, II and $R$ of a typical differential equation-based model), and 2) The presence of probability rates such as $\mu$, $\beta$, $\gamma$ and $\delta$, which implies stochasticity in the timing of birth/death, infection, recovery and immune waning in finite populations.
We implement demographic stochasticity by considering an SIRS model (with vital dynamics) in which individuals are discrete, and may transition between compartments at rates given by (6) (Materials and Methods). Following the approach of (51, 52, 53), based on the van Kampen system size expansion, we derive an expression for the power spectral density (PSD) $P(\omega)$ of noise-induced fluctuations in the (non-seasonal) SIRS model with demographic stochasticity:
$P(\omega)=\frac{\alpha+B\omega^{2}}{(\omega^{2}-\Omega_{0}^{2})^{2}+\Gamma^{2}\omega^{2}}.$ (1)
Here, $\alpha$, $B$, $\Omega_{0}^{2}$, and $\Gamma^{2}$ are all functions of the parameters $\beta$, $\gamma$, $\delta$ and $\mu$. Their expressions are given in the supplement along with a derivation of (1). In 2A, we plot the PSD with parameters from the 2010–2025 fit, and note that while the peak is at five years, a range of frequencies are excited by demographic noise.
By simulating the seasonally-forced stochastic SIRS model using the Doob-Gillespie algorithm (details in Materials and Methods) (54), we find that multi-year cycles may indeed be sustained by demographic noise in sufficiently small populations, suggesting that the system exhibits a stochastic resonance (46). However, the effects of demographic stochasticity are weaker in larger populations, and clear multiannual cycles can only be sustained by this mechanism in populations of less than approximately 2 million individuals, see Fig. 2. Simple demographic stochasticity alone is thus unlikely to be the source of sustained multiannual cycles in M. pneumoniae, at least as far as well-mixed populations are considered. In the supplement, we consider a metapopulation version of the SIRS model with demographic stochasticity and show that multiannual cycles may persist in larger, incompletely mixing populations (see Supplementary Fig. S11).
@@
We go on to consider the effects of extra-demographic or environmental stochasticity. Random fluctuations in the transmission rate $\beta(t)$ will be the primary source of environmental stochasticity considered, but we will also explore the effects of more generic exogenous shocks to the system. Such perturbations can have many different origins, including changes in contact and mobility patterns (such as during holiday periods or due to large events), superspreading of biological and behavioral origin (55, 56, 57, 58), as well as phenotypic variation in the pathogen itself due to mutations. We implement a version of the seasonal SIRS model with stochastic transmission rates based on the method given in (59) (see Materials and Methods). This SIRS model with environmental stochasticity is then used for parameter estimation as well as direct simulation. The parameters obtained by fitting this model to 2010–2025 data can be found in Table 1, labeled “stoch.”.
@@
As shown in Fig. 3, random fluctuations in the transmission rate tend to stimulate approximately 5-year cycles for a wide range of moderate noise levels. When stochasticity is very slight, the annual cycle dominates (e.g. $\sigma=0.05$ in Fig. 3), and when environmental fluctuations are very strong, the epidemic dynamics lacks predictable patterns and becomes dominated by large but rare outbreaks, appearing as a Fourier spectrum with all frequencies represented and no discernible peaks (see e.g. $\sigma=0.5$ in Fig. 3). It should be noted that even a subdominant multiannual peak in the Fourier spectrum (with e.g. half of the peak power of the annual signal) leads to a prevalence time series with a clear multiannual component. The averaging process of Fig. 3 leads to broader and flatter multiannual peaks than those found in individual spectra, but clearly shows the range of excited periodicities.
@@
While the 5-year cycle is the most typical multiannual cycle at moderate noise levels (given M. pneumoniae parameters), other cycles are excited as well, as shown in Fig. 3, with approximately three-year as well as 6–7 year cycles being somewhat common at moderate-to-high noise levels ($\sigma>0.2$). In Supplementary Fig. S4, we show how the addition of continuous environmental noise to an initially purely annual system can rapidly excite the multiannual cycles.